OpenStudy (pinknabastak):

HELLLPPPPPPP!!!!! find the 8th term:

OpenStudy (pinknabastak):

\[(2x-4)^{13}\]

OpenStudy (jdoe0001):

have you covered the "binomial theorem" yet?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=1pSD8cYyqUo

OpenStudy (pinknabastak):

i took the class last year but i never took the final exam so my teacher gave me until spring to finish it. i did go over that theorem but last year so i do not remember..

OpenStudy (anonymous):

\[Tr+1=\left(\begin{matrix}n \\r\end{matrix}\right)y ^{n-r}a ^{r}\in\left( y+a \right)^{n}\] put y=2x,a=-4,n=13,r+1=8

OpenStudy (pinknabastak):

woahhhh....wutt????

OpenStudy (pinknabastak):

@RadEn pleaseee helppp

OpenStudy (raden):

well, the 8th term of (2x -4)^13 is 13C7 (2x)^(13-7) * (-4)^7

OpenStudy (pinknabastak):

what? it makes no sense to me

OpenStudy (raden):

there are 3 cases here : 13C7 = ... ? (2x)^6 = ... ? (-4)^7 = ... ?

OpenStudy (pinknabastak):

whats C?

OpenStudy (raden):

C = combination symbol use this formula to get 13C7 nCr = n!/(r!(n-r)!)

OpenStudy (jdoe0001):

\(\bf (2x-4)^{13}\\ \quad \\ \quad \\ (2x)^{13}+13(2x)^{12}(-4)^1+\square (2x)^{11}(-4)^2+\square (2x)^{10}(-4)^3\\ \quad \\+\square (2x)^{9}(-4)^4+ \square (2x)^8(-4)^5+\square (2x)^7(-4)^6...\) notice as the 1st term exponent DECREASES, the 2nd term exponent INCREASES simultaneously, that's pretty straightforward now to find the coefficients, is the only part that you'd need

OpenStudy (pinknabastak):

ok look fir this question i had an answer written down but im not sure if its right i wrote \[12^{13}=1.069932 \times 10^{14}\]

OpenStudy (jdoe0001):

hmmm... well... if you dunno the binomial theorem... how did you get that?

OpenStudy (jdoe0001):

ohh that's another one?

OpenStudy (pinknabastak):

i dont know i think i asked it before but i dont remember it al all so i wanted to make sure it was right

OpenStudy (jdoe0001):

\(\bf 12^{13}=106993205379072\implies 106993205379072\times 10^1\\ \quad \\ \implies 1.06993205379072\times 10^{14}\) why? well, there are 15 digits, you move the decimal to the left 14 spots, and multiply by a likewise base 10, it'd equate, yes

OpenStudy (pinknabastak):

so it is in fact correct...do you by any chance know how i got it.... lol

OpenStudy (raden):

this site might useful. http://www.mathsisfun.com/algebra/binomial-theorem.html

OpenStudy (jdoe0001):

I wonder that myself =) keep in mind that exercises aren't there for answers to be just transcribed over, they're only useful if you do them, no transcribe material over

OpenStudy (pinknabastak):

I JUST FIGURED OUT HOW I GOT IT!!!!

OpenStudy (jdoe0001):

good

OpenStudy (pinknabastak):

i did.... so if im looking for the 8th term i plugged in 8 for x so i did \[(2(8)-4)^{13}\] so then i got \[12^{13}\] and that is what i put

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

setting x = 8, is not the same as the 8th term in the sequence I said your material of \(\bf 12^{13}=106993205379072\) is correct, you never mentioned it was your 8th term, so, no it's not the 8th term x = 8 and the 8th term aren't the same thing check the video above or the urls posted for the binomial theorem... :) is a rather short video you know

OpenStudy (pinknabastak):

can you just give me the answer.....? haha cuz i really dont get it

OpenStudy (jdoe0001):

it's not easy to get it when not reading the material so... check the video or urls

OpenStudy (pinknabastak):

ok thanks...............